OpenStudy (anonymous):

why does x/2-x equal (-x)/2

OpenStudy (anonymous):

\[\frac{x}{2}-x\text{=}-\frac{x}{2}\]Refer to the LHS. -1x take away +1/2 is equal to (-1/2 x), the quatitiy on the RHS.

OpenStudy (anonymous):

Disregard the prior posting. Refer to the LHS. -1x "added to" +1/2x is equal to (-1/2 x), the quantity on the RHS.

OpenStudy (anonymous):

Disregard the prior posting. Refer to the LHS. -1x "added to" +1/2x is equal to (-1/2 x), the quantity on the RHS.

OpenStudy (anonymous):

where is x/2 in the solution you wrote I don't get it sorry

OpenStudy (anonymous):

\[+\frac{x}{2}-x\text{ = }-\frac{x}{2}\]\[-\frac{x}{2}\text{ = }-\frac{x}{2}\]